OpenStudy (anonymous):

Evaluate. Write in simplest form. 6-b / 3a for a = 10, b = –9 A.1/10 B.-3/10 C.2 D.1/2

OpenStudy (hoblos):

(6+9)/30 = 15/30 = 1/2